OpenStudy (anonymous):

Solve the equation Twenty-five minus four b equals negative six b plus forty-seven for b. b equals negative eleven b equals negative two and two tenths b equals two and two tenths b equals eleven

OpenStudy (goldphenoix):

In standard notation, it would look like: \[\large \large 25 - 4b = -6b + 47\] What do you think you do first?

OpenStudy (anonymous):

Subtract

OpenStudy (goldphenoix):

Good. Subtract what?

OpenStudy (anonymous):

25-4?

OpenStudy (goldphenoix):

Hmm? 25 - 4b?

OpenStudy (anonymous):

Yes but i just dont know wat the b stands for

OpenStudy (goldphenoix):

Well. We want to fin db. So we must bring the variable on one side, and the number on the other side. So what can we subtract?

OpenStudy (anonymous):

25-b4?

OpenStudy (goldphenoix):

Nope. I'll give you a headstart. You want to subtract 25 on both side. So what does that give you?

OpenStudy (anonymous):

0

OpenStudy (goldphenoix):

47-25= ?

OpenStudy (anonymous):

22

OpenStudy (goldphenoix):

Good. And 25 - 25 = 0. So your equation should look like: \[\large \large -4b =-6b + 22\]

OpenStudy (goldphenoix):

Now, what do you think you do next?

OpenStudy (anonymous):

Add

OpenStudy (goldphenoix):

Add what?

OpenStudy (anonymous):

46+22

OpenStudy (goldphenoix):

Where did you get 46 from?

OpenStudy (goldphenoix):

Tip: We want to bring -6b to the left. How can we do that?

OpenStudy (anonymous):

I think by putting 4b to the right

OpenStudy (goldphenoix):

We want to do something to both side that has the same value. If we did something to both side that wasn't equal, then it wouldn't be an equation. To bring -6b to the left, you want to add both side by 6b. What do you get?

OpenStudy (goldphenoix):

-6b + 6b = ?

OpenStudy (anonymous):

Ok but the problem im having id the letters like i dont know how to add or subtract the letters with the numbers

OpenStudy (anonymous):

is*

OpenStudy (goldphenoix):

What do you mean? Give me an example. Like 10x + 2x = ?

OpenStudy (anonymous):

Like n=-1

OpenStudy (goldphenoix):

If it's like n = -1. Then you can't do anything more. Since you already find your variable. Do you mean the original problem you had?

OpenStudy (anonymous):

yess

OpenStudy (goldphenoix):

Well. You want to bring the variable on one side, and the number on the other side. So if you have for example: 10 + 2x = 4x + 5 You would want to bring the variable on one side. So let say we want to bring 2x to the right. We would want to subtract both side by 2x . 10 + 2x - 2x = 10. 4x - 2x + 5 = 2x + 5 Does it make sense? If my teaching doesn't make sense, then click here: http://www.onlinemathlearning.com/equations-variables-both-sides.html

OpenStudy (goldphenoix):

 ^Example on how you should solve that example. Make sense?